OpenStudy (brodybarnes35):

For f(x) = x2 and g(x) = (x − 5)2, in which direction and by how many units should f(x) be shifted to obtain g(x)? Up 5 units Down 5 units Left 5 units Right 5 units SOMEONE HELP FAN AND MEDAL

OpenStudy (anonymous):

okay so is that x squared

OpenStudy (anonymous):

or just 2x

OpenStudy (brodybarnes35):

yeah squared

OpenStudy (anonymous):

okay

OpenStudy (brodybarnes35):

@abb0t @BAdhi @Cookie_2046 @dan815 @e.mccormick @Frostbite @ganeshie8 @Hero @inkyvoyd @jim_thompson5910 @keshaun455 @LunyMoony @mathway @Nnesha @oldrin.bataku @pooja195 @QueenofTamaran @Robert136 @sammixboo @tHe_FiZiCx99 @uybuyvf @Vocaloid @welshfella @xapproachesinfinity

OpenStudy (anonymous):

okay, so we have \[x ^{2}\] and \[x ^{2}-10x+25\]

OpenStudy (brodybarnes35):

ok so what does that mean?

OpenStudy (anonymous):

from your answer choices, how are you going to get the first part to the second

OpenStudy (brodybarnes35):

idk, idk how to do this lol

OpenStudy (anonymous):

wait, you what would be easier? lol scratch everything i said

OpenStudy (anonymous):

so

OpenStudy (brodybarnes35):

ok :)

OpenStudy (anonymous):

since both expressions are squared, we can square root both of them

OpenStudy (xapproachesinfinity):

you better live it g(x)=(x-5)^2 to see what shift needed

OpenStudy (anonymous):

meaning we just have x and x-5

OpenStudy (xapproachesinfinity):

leave*

OpenStudy (anonymous):

so, since you are subtracting, you are either going down or to the left

OpenStudy (brodybarnes35):

ok

OpenStudy (anonymous):

i wouldnt leave it if i were you actually

OpenStudy (brodybarnes35):

ok go on please :)

OpenStudy (xapproachesinfinity):

ok let's how you gonna do it

OpenStudy (xapproachesinfinity):

see*

OpenStudy (brodybarnes35):

ok

OpenStudy (anonymous):

well we already have x and x-5, because we square rooted, and you want him to leave it as squared, thats totally fine and you can help him

OpenStudy (xapproachesinfinity):

that's not a square root? where did you see any square root?

OpenStudy (xapproachesinfinity):

Okay lets consider a function f(x)=x^2 we call it parent function

OpenStudy (brodybarnes35):

ok

OpenStudy (xapproachesinfinity):

if we want to move the graph of that function f(x) to right some units away i will use b units here we just do the following \(f(x-b)=(x-b)^2\) that is b units to the right away ffrom f(x) let's give this new function a name h(x) so that \(h(x)=f(x-b)=(x-b)^2\)

OpenStudy (xapproachesinfinity):

following ?

OpenStudy (xapproachesinfinity):

now we want to move f(x) to obtain g(x) look at the g(x) and tell me what do you need to do to move f such that it becomes g

OpenStudy (xapproachesinfinity):

earth to you, hello, are you there?

OpenStudy (brodybarnes35):

yes im following

OpenStudy (xapproachesinfinity):

so according to what i just said what can we do to shift f to get g?

OpenStudy (xapproachesinfinity):

just compare what i did with what we have?

OpenStudy (brodybarnes35):

i honestly dont know im guessing 5 down or to the left

OpenStudy (xapproachesinfinity):

why down and left in my explanation i have not spelled at all any thing about down or left shifting

OpenStudy (xapproachesinfinity):

i only i explained one thing; that is shift a curve x^2 to right by a b units to get (x-b)^2

OpenStudy (xapproachesinfinity):

so basically to get from f(x)=x^2 to g(x)=(x-5)^2 we need to shift f(x) 5 units to the right

OpenStudy (brodybarnes35):

can you just tell me the answer and explain cause you're not helping that well :)

OpenStudy (brodybarnes35):

oh thanks

OpenStudy (brodybarnes35):

so d?

OpenStudy (xapproachesinfinity):

it is not me who is not helping it is you who don't want to be helped haha

OpenStudy (xapproachesinfinity):

i was deriving you to the answer but you seem not to pay attention

OpenStudy (xapproachesinfinity):

yes it is D

OpenStudy (brodybarnes35):

ok thanks can you help on like 4 more???

OpenStudy (brodybarnes35):

are u sure?

OpenStudy (brodybarnes35):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

trillion times sure! and no i can't if you don't have much interest to work the problem

OpenStudy (brodybarnes35):

no i will i promise @xapproachesinfinity

OpenStudy (brodybarnes35):

What is the domain of the following parabola? u-shaped graph that opens up with a vertex of 2, negative 3 All real numbers x ≥ 1 x ≤ 2 y ≥ −3

OpenStudy (brodybarnes35):

@xapproachesinfinity

OpenStudy (brodybarnes35):

help me @xapproachesinfinity

OpenStudy (xapproachesinfinity):

all sort of parabolas have the domain all real number no matter what it is

OpenStudy (xapproachesinfinity):

if we speak about co-domain (range) then the story is different

OpenStudy (brodybarnes35):

so its all real numbers im assuming?wait no

OpenStudy (xapproachesinfinity):

if you want the reason just think is there any number that cause a function of the sort f(x)=ax^2+bx+c does not exist i mean is there a point where that graph has discontinuity

OpenStudy (xapproachesinfinity):

the answer is no, all real numbers are acceptable in the machine of parabolas

OpenStudy (brodybarnes35):

so no all real numbers

OpenStudy (xapproachesinfinity):

im saying all real numbers and you are saying no all real numbers

OpenStudy (xapproachesinfinity):

pay a bit of attention dear :)

OpenStudy (brodybarnes35):

ohhh ok it is all real numbers i apologize

OpenStudy (xapproachesinfinity):

no problem

OpenStudy (brodybarnes35):

The quadratic functions f(x) and g(x) are described as follows: f(x) = −8x2 + 7 x g(x) 0 0 1 2 2 6 3 2 4 0 Which of the following statements best compares the maximum value of the 2 functions? It is the same for both functions. f(x) has a greater maximum value than g(x). g(x) has a greater maximum value than f(x). The maximum values cannot be determined can u help on like 3 more? plz @xapproachesinfinity

OpenStudy (brodybarnes35):

@xapproachesinfinity

OpenStudy (brodybarnes35):

@abb0t @BAdhi @Cookie_2046 @dan815 @e.mccormick @Frostbite @ganeshie8 @Hero @inkyvoyd @jim_thompson5910 @keshaun455 @LunyMoony @mathway @Nnesha @oldrin.bataku @pooja195 @QueenofTamaran @Robert136 @sammixboo @tHe_FiZiCx99 @uybuyvf @Vocaloid @welshfella @xapproachesinfinity

OpenStudy (brodybarnes35):

@Sammylemon